# Flexible Deep Neural Networks for Partially Linear Survival Data

### v1

@@ 8 Appendix - Proofs
Suppose Assumptions A1A2 and A4 hold. Let $g_{0},g_{1}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ and fix $\mbox{${\theta}$}\in\Theta$. For $j=0,1$, let $P_{g_{j}}$ denote the distribution of $(T,{\bf X},{\bf Z},\Delta)$ under event hazard $h_{g_{j}}(t\mid{\bf X},{\bf Z})=\exp\{g_{j}(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}$ and a common censoring hazard $\mu(t\mid{\bf X},{\bf Z})$. As a consequence of Assumptions A1A2, there exists a constant $0<D<\infty$ such that, almost surely,
@@
Define $c=e^{-D}$ and $C=e^{D}$. Then, for any nonnegative measurable function $q:[0,\tau]\times\mathcal{X}\to[0,\infty)$ with $\mathbb{E}_{P_{g_{0}}}[q(T,{\bf X})]<\infty$,
$K_{-}\,\mathbb{E}_{P_{g_{0}}}\!\big\{q(T,{\bf X})\big\}\;\leq\;\mathbb{E}_{P_{g_{1}}}\!\big\{q(T,{\bf X})\big\}\;\leq\;K_{+}\,\mathbb{E}_{P_{g_{0}}}\!\big\{q(T,{\bf X})\big\}\,$,
@@
In particular, $\mathbb{E}_{P_{g_{1}}}[q(T,{\bf X})]\asymp\mathbb{E}_{P_{g_{0}}}[q(T,{\bf X})]$ with constants depending only on $(c,C,\tau)$.
@@
$K_{-}\;\leq\;\frac{dP_{g_{1}}}{dP_{g_{0}}}(T,\Delta,{\bf X},{\bf Z})\;\leq\;K_{+},\qquad P_{g_{0}}\text{--a.s.},$
where $K_{-}=\frac{c}{C}e^{-\tau(C-c)}$ and $K_{+}=\frac{C}{c}e^{\tau(C-c)}$. For any nonnegative measurable function $q$ with finite expectation under $P_{g_{0}}$,
$\mathbb{E}_{P_{g_{1}}}\{q(T,{\bf X})\}=\mathbb{E}_{P_{g_{0}}}\!\left\{q(T,{\bf X})\,\frac{dP_{g_{1}}}{dP_{g_{0}}}(T,\Delta,{\bf X},{\bf Z})\right\}$
@@
$K_{-}\,\mathbb{E}_{P_{g_{0}}}\{q(T,{\bf X})\}\;\leq\;\mathbb{E}_{P_{g_{1}}}\{q(T,{\bf X})\}\;\leq\;K_{+}\,\mathbb{E}_{P_{g_{0}}}\{q(T,{\bf X})\}\,.$
@@
Assume A2. Let $\eta_{j}=(\mbox{${\theta}$}_{j},g_{j})$, $j=1,2$, be two parameter values in $(\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M))\cup(\Theta\times\mathcal{G}(K,s,\boldsymbol{p},D))$. Define
@@
If Assumptions A1A5 hold, then there exists a constant $c_{\theta}>0$ depending only on the constants in A1A4 such that, for all parameters in a neighborhood of $\eta_{o}$,
@@
Let $\mbox{${\beta}$}=\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}$, $a(t,{\bf X})=g_{1}(t,{\bf X})-g_{2}(t,{\bf X})$, $b=\mbox{${\beta}$}^{\top}{\bf Z}$. At a failure time, $\chi_{\eta_{1}}-\chi_{\eta_{2}}=a+b$. Thus
$d(\eta_{1},\eta_{2})^{2}=\mathbb{E}_{0}\{\Delta(a+b)^{2}\}\,.$
To find the smallest possible value of this expression over all possible nuisance differences $g_{1}-g_{2}$, minimize $(a+b)^{2}$ over scalar functions $a(T,{\bf X})$. The minimizing choice is
$a^{\ast}(T,{\bf X})=-\mathbb{E}_{0}\{b\mid T,{\bf X},\Delta=1\}\,.$
Substituting this optimal choice yields
$\inf_{g_{1}-g_{2}}d(\eta_{1},\eta_{2})^{2}=\mathbb{E}_{0}\!\Big\{\Delta\,Var(b\mid T,{\bf X},\Delta=1)\Big\}\,.$
This is exactly the standard least–squares argument. Assumption A4 ensures
$\Pr(\Delta=1\mid T,{\bf X})\;\geq\;c_{\Delta}\quad\text{a.s.}$
@@
$\mathbb{E}_{0}\Big[\Pr(\Delta=1\mid T,{\bf X})\,Var(b\mid T,{\bf X},\Delta=1)\Big]\;\geq\;c_{\Delta}\,\mathbb{E}_{0}\bigl[Var(b\mid T,{\bf X},\Delta=1)\bigr]\,.$
Assumption A5 implies that for some $c_{Z}>0$,
$Var({\bf Z}\mid T,{\bf X},\Delta=1)\succeq c_{Z}I_{p}\,$,
thus for any $\beta$},
$Var(b\mid T,{\bf X},\Delta=1)=Var(\mbox{${\beta}$}^{\top}{\bf Z}\mid T,{\bf X},\Delta=1)\geq c_{\Delta}c_{Z}\|\mbox{${\beta}$}\|^{2}\,.$
Combining the above, we get
$d(\eta_{1},\eta_{2})^{2}\geq c_{Z}\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}.$
Setting $c_{\theta}=c_{\Delta}c_{Z}>0$ completes the proof. ∎
@@
$\mathbb{E}^{*}\!\sup_{d(\eta,\eta_{o})\leq\delta}\bigl||\mathbb{G}_{n}\big\{\ell(\eta;\cdot)-\ell(\eta_{o};\cdot)\big\}\bigr|\;\leq\;\Psi_{n}(\delta)\equiv\delta\,\sqrt{s\log\frac{\mathcal{U}}{\delta}}+\frac{s}{\sqrt{n}}\log\frac{\mathcal{U}}{\delta}\,$, (12)
@@
Set $\tau_{n}=\gamma_{n}\log^{2}n$ as in the theorem. By Assumption A3 and the specific choice of $(K,s,\boldsymbol{p})$ (cf. the construction in schmidthieber2020nonparametric and zhong2022deep), we have
@@
$\Psi_{n}(\tau_{n})\;\leq\;\frac{c}{2}\,\sqrt{n}\,\tau_{n}^{2}\quad\mbox{and}\quad\frac{1}{\sqrt{n}}\Psi_{n}(\tau_{n})\;\leq\;\frac{c}{2}\,\tau_{n}^{2}\,.$ (13)
@@
$\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|(\ell_{n}-\ell_{0})(\eta)-(\ell_{n}-\ell_{0})(\eta_{o})\bigr|=\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\frac{1}{\sqrt{n}}\bigl|\mathbb{G}_{n}\big\{\ell(\eta;\cdot)-\ell(\eta_{o};\cdot)\big\}\bigr|=O_{p}\bigl(\tfrac{1}{\sqrt{n}}\Psi_{n}(\tau_{n})\bigr)=O_{p}(\tau_{n}^{2}),$ (14)
where we used (12) and (13).
By Theorem 1 of schmidthieber2020nonparametric, applied to the composite Hölder class $\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{\bf d},M)$ with the choice of $(K,s,\boldsymbol{p},D)$ given in Assumption A3, there exists $\tilde{g}_{1}\in\mathcal{G}(K,s,\boldsymbol{p},D)$ such that
@@
Applying this to $\eta=\tilde{\eta}_{1}$ and using (15), we obtain
$\ell_{0}(\tilde{\eta}_{1})-\ell_{0}(\eta_{o})\;\geq\;-\,c_{+}\,d(\tilde{\eta}_{1},\eta_{o})^{2}\;\geq\;-\,c_{+}C_{1}\,\tau_{n}^{2}\;=\;-\,C_{2}\,\tau_{n}^{2}\,$, (16)
@@
The empirical-process bound (14) and the likelihood chain (18) verify the conditions (i)–(ii) of Theorem 3.4.1 in vdvwellner1996 with radius of order $\tau_{n}$. Therefore,
@@
$\mathbb{E}\Bigl[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\Bigr]\;\leq\;C\Bigl(d(\widehat{\eta},\eta_{o})^{2}+\|\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|^{2}\Bigr)\,.$ (19)
@@
Hence, $\|\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|=O_{p}(\tau_{n})$, and the right-hand side of (19) is $O_{p}(\tau_{n}^{2})$, which is exactly the claim of Theorem 1. The expectation is taken with respect to a new observation $(T,\Delta,{\bf X},{\bf Z})\sim P_{0}$, while the $O_{p}(\cdot)$ is with respect to the randomness in the training sample used to construct $\widehat{g}$. ∎
@@
For $g\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ and and $\mbox{${\theta}$}_{o}\in\Theta$, let $P_{g,\theta_{o}}$ denote the joint distribution of $(T,\Delta,{\bf X},{\bf Z})$ induced by the hazard $h_{g,\theta_{o}}(t\mid{\bf X},{\bf Z})=\exp\{g(t,{\bf X})+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\}$ together with a censoring hazard that is held fixed across all $g$. By the standard Kullback-Leibler (KL) distance for bounded hazards (Assumptions A1A2), there exists $C>0$ such that for any $g_{0},g_{1}\in\mathcal{\mathcal{}}{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$,
$\mbox{KL}(P_{g_{1},\theta_{o}}\,\|\,P_{g_{0},\theta_{o}})\;\leq\;C\,n\,\mathbb{E}_{P_{g_{1},\theta_{o}}}\!\left[\Delta\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\right]\,.$
@@
$\mbox{KL}(P_{g_{1},\theta_{o}}\,\|\,P_{g_{0},\theta_{o}})\;\leq\;C\,n\,\mathbb{E}_{P_{g_{1},\theta_{o}}}\!\bigl[\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\bigr].$
@@
$\mbox{KL}(P_{g_{1},\theta_{o}}\,\|\,P_{g_{0},\theta_{o}})\;\leq\;C^{\prime}\,n\,\mathbb{E}_{P_{g_{0},\theta_{o}}}\!\bigl[\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\bigr]$
@@
By the packing construction for composite Hölder networks (schmidthieber2020nonparametric), there exist functions $g^{(0)},\dots,g^{(N)}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ and constants $c_{1},c_{2}>0$ such that
$\|g^{(j)}-g^{(k)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}\;\geq\;2c_{1}\,\gamma_{n}\,,\qquad j\neq k\,$, (20)
@@
$\frac{1}{N}\sum_{j=1}^{N}\|g^{(j)}-g^{(0)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}^{2}\;\leq\;\frac{c_{2}\,\log N}{n}\,.$ (21)
Applying the KL bound with $g_{0}=g^{(0)}$ and $g_{1}=g^{(j)}$ and the using (21),
$\frac{1}{N}\sum_{j=1}^{N}\mbox{KL}\!\big(P_{g^{(j)},\theta_{o}}\,\|\,P_{g^{(0)},\theta_{o}}\big)\;\leq\;C^{\prime}\,c_{2}\,\log N\,.$
@@
$\frac{1}{N}\sum_{j=1}^{N}\mbox{KL}(P_{g^{(j)},\theta_{o}}\,\|\,P_{g^{(0)},\theta_{o}})\;\leq\;\alpha\log N\,.$
Applying Tsybakov’s Fano inequality to $\{g^{(0)},\dots,g^{(N)}\}$ with separation (20), we obtain that there exists $c>0$ (depending only on $c_{1}$ and $\alpha$) such that for any estimator $\widehat{g}$,
$\sup_{j=0,\dots,N}\ \mathbb{E}_{P_{g^{(j)},\theta_{o}}}\!\Big(\|\widehat{g}-g^{(j)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}^{2}\Big)\;\geq\;c\,\gamma_{n}^{2}\,.$
@@
$\inf_{\widehat{g}}\ \sup_{g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\Bigl(\|\widehat{g}-g_{o}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}^{2}\Bigr)\;\geq\;c\,\gamma_{n}^{2}\,.$
@@
$\mathbb{E}_{P_{g^{(0)},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\leq\;C_{\mathrm{eq}}\,\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big].$
@@
$\inf_{\widehat{g}}\ \sup_{g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;\frac{c}{C_{\mathrm{eq}}}\,\gamma_{n}^{2}.$
@@
$\inf_{\widehat{g}}\ \sup_{\theta_{o}\in\Theta,\ g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;c_{1}\,\gamma_{n}^{2}$ (22)
@@
$\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\leq\;C_{\Delta}\,\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\,.$
Combining this with (22) gives
$\inf_{\widehat{g}}\ \sup_{\theta_{o}\in\Theta,\ g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;\frac{c_{1}}{C_{\Delta}}\,\gamma_{n}^{2}\,.$

### v2

@@ Appendix A Proofs for Section 4
Suppose Assumptions A1A2 and A4 hold. Let $g_{0},g_{1}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$ and fix $\mbox{${\theta}$}\in\Theta$. For $j=0,1$, let $\mathbb{P}_{g_{j}}$ denote the distribution of $(T,{\bf X},{\bf Z},\Delta)$ under event hazard $h_{g_{j}}(t\mid{\bf X},{\bf Z})=\exp\{g_{j}(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}$ and a common censoring hazard $\mu(t\mid{\bf X},{\bf Z})$. As a consequence of Assumptions A1A2, there exists a constant $0<D<\infty$ such that, almost surely,
@@
Define $c=e^{-D}$ and $C=e^{D}$. Then, for any nonnegative measurable function $q:[0,\tau]\times\mathcal{X}\to[0,\infty)$ with $\mathbb{E}_{\mathbb{P}_{g_{0}}}[q(T,{\bf X})]<\infty$,
$K_{-}\,\mathbb{E}_{\mathbb{P}_{g_{0}}}\!\big\{q(T,{\bf X})\big\}\;\leq\;\mathbb{E}_{\mathbb{P}_{g_{1}}}\!\big\{q(T,{\bf X})\big\}\;\leq\;K_{+}\,\mathbb{E}_{\mathbb{P}_{g_{0}}}\!\big\{q(T,{\bf X})\big\}\,$,
@@
In particular, $\mathbb{E}_{\mathbb{P}_{g_{1}}}[q(T,{\bf X})]\asymp\mathbb{E}_{\mathbb{P}_{g_{0}}}[q(T,{\bf X})]$ with constants depending only on $(c,C,\tau)$.
@@
$K_{-}\;\leq\;\frac{d\mathbb{P}_{g_{1}}}{d\mathbb{P}_{g_{0}}}(T,\Delta,{\bf X},{\bf Z})\;\leq\;K_{+},\qquad\mathbb{P}_{g_{0}}\text{--a.s.},$
where $K_{-}=\frac{c}{C}e^{-\tau(C-c)}$ and $K_{+}=\frac{C}{c}e^{\tau(C-c)}$. For any nonnegative measurable function $q$ with finite expectation under $\mathbb{P}_{g_{0}}$,
$\mathbb{E}_{\mathbb{P}_{g_{1}}}\{q(T,{\bf X})\}=\mathbb{E}_{\mathbb{P}_{g_{0}}}\!\left\{q(T,{\bf X})\,\frac{d\mathbb{P}_{g_{1}}}{d\mathbb{P}_{g_{0}}}(T,\Delta,{\bf X},{\bf Z})\right\}$
@@
$K_{-}\,\mathbb{E}_{\mathbb{P}_{g_{0}}}\{q(T,{\bf X})\}\;\leq\;\mathbb{E}_{\mathbb{P}_{g_{1}}}\{q(T,{\bf X})\}\;\leq\;K_{+}\,\mathbb{E}_{\mathbb{P}_{g_{0}}}\{q(T,{\bf X})\}\,.$
@@
Assume A2. Let $\eta_{j}=(\mbox{${\theta}$}_{j},g_{j})$, $j=1,2$, be two parameter values in $(\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M))\cup(\Theta\times\mathcal{G}(K,s,\boldsymbol{p},D))$. Define
@@
If Assumptions A1A5 hold, then there exists a constant $c_{\theta}>0$ such that, for all parameters $\eta_{1}=(\mbox{${\theta}$}_{1},g_{1})$ and $\eta_{2}=(\mbox{${\theta}$}_{2},g_{2})$,
@@
Let $\mbox{${\beta}$}=\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}$, $a(T,{\bf X})=g_{1}(T,{\bf X})-g_{2}(T,{\bf X})$, $b=\mbox{${\beta}$}^{\top}{\bf Z}$. Then
$\chi_{\eta_{1}}(T,{\bf X},{\bf Z})-\chi_{\eta_{2}}(T,{\bf X},{\bf Z})=a(T,{\bf X})+b,$
and hence
$d(\eta_{1},\eta_{2})^{2}=\mathbb{E}_{0}\!\left[\Delta\{a(T,{\bf X})+b\}^{2}\right].$
Because $a(T,{\bf X})$ is measurable with respect to $(T,{\bf X})$, we have
$\mathbb{E}_{0}\!\left[\Delta\{a+b\}^{2}\mid T,{\bf X}\right]=\Pr(\Delta=1\mid T,{\bf X})\,\mathbb{E}_{0}\!\left[\{a+b\}^{2}\mid T,{\bf X},\Delta=1\right].$
By the conditional variance decomposition,
$\mathbb{E}_{0}\!\left[\{a+b\}^{2}\mid T,{\bf X},\Delta=1\right]=\Big(a+\mathbb{E}_{0}[b\mid T,{\bf X},\Delta=1]\Big)^{2}+Var_{0}(b\mid T,{\bf X},\Delta=1).$
@@
$\mathbb{E}_{0}\!\left[\Delta\{a+b\}^{2}\right]=\mathbb{E}_{0}\!\left[\Pr(\Delta=1\mid T,{\bf X})\Big(a+\mathbb{E}_{0}[b\mid T,{\bf X},\Delta=1]\Big)^{2}\right]+\mathbb{E}_{0}\!\left[\Delta\,Var_{0}(b\mid T,{\bf X},\Delta=1)\right].$
Since the first term is nonnegative,
$d(\eta_{1},\eta_{2})^{2}\geq\mathbb{E}_{0}\!\left[\Delta\,Var_{0}(b\mid T,{\bf X},\Delta=1)\right].$
Now, because $b=\mbox{${\beta}$}^{\top}{\bf Z}$,
$Var_{0}(b\mid T,{\bf X},\Delta=1)=\mbox{${\beta}$}^{\top}Var_{0}({\bf Z}\mid T,{\bf X},\Delta=1)\mbox{${\beta}$},$
and thus
$\mathbb{E}_{0}\!\left[\Delta\,Var_{0}(b\mid T,{\bf X},\Delta=1)\right]=\mbox{${\beta}$}^{\top}\mathbb{E}_{0}\!\left[\Delta\,\Big\{{\bf Z}-\mathbb{E}_{0}({\bf Z}\mid T,{\bf X},\Delta=1)\Big\}^{\otimes 2}\right]\mbox{${\beta}$}=\mbox{${\beta}$}^{\top}\Sigma_{{\bf Z}}\mbox{${\beta}$}.$
By Assumption A5,
$\mbox{${\beta}$}^{\top}\Sigma_{{\bf Z}}\mbox{${\beta}$}\geq\lambda_{\min}(\Sigma_{\bf Z})\,\|\mbox{${\beta}$}\|^{2}.$
Hence
$d(\eta_{1},\eta_{2})^{2}\geq\lambda_{\min}(\Sigma_{\bf Z})\,\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}.$
This proves the lemma with $c_{\theta}=\lambda_{\min}(\Sigma_{\bf Z})$. ∎
@@
$\mathbb{E}^{*}\!\sup_{d(\eta,\eta_{o})\leq\delta}\bigl|\mathbb{G}_{n}\big\{\ell(\eta;\cdot)-\ell(\eta_{o};\cdot)\big\}\bigr|\;\leq\;\Psi_{n}(\delta)\equiv\delta\,\sqrt{s\log\frac{\mathcal{U}}{\delta}}+\frac{s}{\sqrt{n}}\log\frac{\mathcal{U}}{\delta}\,$, (A.3)
@@
Set $\tau_{n}=\gamma_{n}\log^{2}n$ as in the theorem. By Assumption A3 and the specific choice of $(K,s,\boldsymbol{p})$ (cf. the construction in Schmidt-Hieber (2020) and Zhong et al. (2022)), we have
@@
$\Psi_{n}(\tau_{n})\;\leq\;\frac{c}{2}\,\sqrt{n}\,\tau_{n}^{2}\quad\mbox{and}\quad\frac{1}{\sqrt{n}}\Psi_{n}(\tau_{n})\;\leq\;\frac{c}{2}\,\tau_{n}^{2}\,.$ (A.4)
@@
$\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|(\ell_{n}-\ell_{0})(\eta)-(\ell_{n}-\ell_{0})(\eta_{o})\bigr|=\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\frac{1}{\sqrt{n}}\bigl|\mathbb{G}_{n}\big\{\ell(\eta;\cdot)-\ell(\eta_{o};\cdot)\big\}\bigr|=O_{p}\bigl(\tfrac{1}{\sqrt{n}}\Psi_{n}(\tau_{n})\bigr)=O_{p}(\tau_{n}^{2}),$ (A.5)
where we used (A.3) and (A.4).
By Theorem 1 of Schmidt-Hieber (2020), applied to the composite Hölder class $\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{\bf d},M)$ with the choice of $(K,s,\boldsymbol{p},D)$ given in Assumption A3, there exists $\tilde{g}_{1}\in\mathcal{G}(K,s,\boldsymbol{p},D)$ such that
@@
Applying this to $\eta=\tilde{\eta}_{1}$ and using (A.6), we obtain
$\ell_{0}(\tilde{\eta}_{1})-\ell_{0}(\eta_{o})\;\geq\;-\,c_{+}\,d(\tilde{\eta}_{1},\eta_{o})^{2}\;\geq\;-\,c_{+}C_{1}\,\tau_{n}^{2}\;=\;-\,C_{2}\,\tau_{n}^{2}\,$, (A.7)
@@
The empirical-process bound (A.5) and the likelihood chain (A.9) verify the conditions (i)–(ii) of Theorem 3.4.1 in van der Vaart and Wellner (1996) with radius of order $\tau_{n}$. Therefore,
@@
$\mathbb{E}\Bigl[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\Bigr]\;\leq\;C\Bigl(d(\widehat{\eta},\eta_{o})^{2}+\|\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|^{2}\Bigr)\,.$ (A.10)
@@
Hence, $\|\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|=O_{p}(\tau_{n})$, and the right-hand side of (A.10) is $O_{p}(\tau_{n}^{2})$, which is exactly the claim of Theorem 1. The expectation is taken with respect to a new observation $(T,\Delta,{\bf X},{\bf Z})\sim\mathbb{P}_{0}$, while the $O_{p}(\cdot)$ is with respect to the randomness in the training sample used to construct $\widehat{g}$. ∎
@@
For $g\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$ and $\mbox{${\theta}$}_{o}\in\Theta$, let $\mathbb{P}_{g,\theta_{o}}$ denote the joint distribution of $(T,\Delta,{\bf X},{\bf Z})$ induced by the hazard $h_{g,\theta_{o}}(t\mid{\bf X},{\bf Z})=\exp\{g(t,{\bf X})+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\}$ together with a censoring hazard that is held fixed across all $g$. By the standard Kullback-Leibler (KL) distance for bounded hazards (Assumptions A1A2), there exists $C>0$ such that for any $g_{0},g_{1}\in\mathcal{\mathcal{}}{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$,
$\mbox{KL}(\mathbb{P}_{g_{1},\theta_{o}}\,\|\,\mathbb{P}_{g_{0},\theta_{o}})\;\leq\;C\,n\,\mathbb{E}_{\mathbb{P}_{g_{1},\theta_{o}}}\!\left[\Delta\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\right]\,.$
@@
$\mbox{KL}(\mathbb{P}_{g_{1},\theta_{o}}\,\|\,\mathbb{P}_{g_{0},\theta_{o}})\;\leq\;C\,n\,\mathbb{E}_{\mathbb{P}_{g_{1},\theta_{o}}}\!\bigl[\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\bigr].$
@@
$\mbox{KL}(\mathbb{P}_{g_{1},\theta_{o}}\,\|\,\mathbb{P}_{g_{0},\theta_{o}})\;\leq\;C^{\prime}\,n\,\mathbb{E}_{\mathbb{P}_{g_{0},\theta_{o}}}\!\bigl[\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\bigr]$
@@
By the packing construction for composite Hölder networks (Schmidt-Hieber, 2020), there exist functions $g^{(0)},\dots,g^{(N)}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$ and constants $c_{1},c_{2}>0$ such that
$\|g^{(j)}-g^{(k)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}\;\geq\;2c_{1}\,\gamma_{n}\,,\qquad j\neq k\,$, (A.11)
@@
$\frac{1}{N}\sum_{j=1}^{N}\|g^{(j)}-g^{(0)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}^{2}\;\leq\;\frac{c_{2}\,\log N}{n}\,.$ (A.12)
Applying the KL bound with $g_{0}=g^{(0)}$ and $g_{1}=g^{(j)}$ and using (A.12),
$\frac{1}{N}\sum_{j=1}^{N}\mbox{KL}\!\big(\mathbb{P}_{g^{(j)},\theta_{o}}\,\|\,\mathbb{P}_{g^{(0)},\theta_{o}}\big)\;\leq\;C^{\prime}\,c_{2}\,\log N\,.$
@@
$\frac{1}{N}\sum_{j=1}^{N}\mbox{KL}(\mathbb{P}_{g^{(j)},\theta_{o}}\,\|\,\mathbb{P}_{g^{(0)},\theta_{o}})\;\leq\;\alpha\log N\,.$
Applying Tsybakov’s Fano inequality to $\{g^{(0)},\dots,g^{(N)}\}$ with separation (A.11), we obtain that there exists $c>0$ (depending only on $c_{1}$ and $\alpha$) such that for any estimator $\widehat{g}$,
$\sup_{j=0,\dots,N}\ \mathbb{E}_{\mathbb{P}_{g^{(j)},\theta_{o}}}\!\Big(\|\widehat{g}-g^{(j)}\|_{L^{2}(\mathbb{P}_{g^{(0)},\theta_{o}})}^{2}\Big)\;\geq\;c\,\gamma_{n}^{2}\,.$
@@
$\inf_{\widehat{g}}\ \sup_{g_{o}\in\mathcal{H}}\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\Bigl(\|\widehat{g}-g_{o}\|_{L^{2}(\mathbb{P}_{g^{(0)},\theta_{o}})}^{2}\Bigr)\;\geq\;c\,\gamma_{n}^{2}\,.$
@@
$\mathbb{E}_{\mathbb{P}_{g^{(0)},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\leq\;C_{\mathrm{eq}}\,\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big].$
@@
$\inf_{\widehat{g}}\ \sup_{g_{o}\in\mathcal{H}}\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;\frac{c}{C_{\mathrm{eq}}}\,\gamma_{n}^{2}.$
@@
$\inf_{\widehat{g}}\ \sup_{\theta_{o}\in\Theta,\ g_{o}\in\mathcal{H}}\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;c_{1}\,\gamma_{n}^{2}$ (A.13)
@@
$\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\leq\;C_{\Delta}\,\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\big[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\,.$
Combining this with (A.13) gives
$\inf_{\widehat{g}}\ \sup_{\theta_{o}\in\Theta,\ g_{o}\in\mathcal{H}}\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\big[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;\frac{c_{1}}{C_{\Delta}}\,\gamma_{n}^{2}\,.$